# Video Soundtrack Generation by Aligning Emotions and Temporal Boundaries

### v2

@@ Abstract
We introduce EMSYNC, a video-based symbolic music generation model that aligns music with a video’s emotional content and temporal boundaries. It follows a two-stage framework, where a pretrained video emotion classifier extracts emotional features, and a conditional music generator produces MIDI sequences guided by both emotional and temporal cues. We introduce boundary offsets, a novel temporal conditioning mechanism that enables the model to anticipate upcoming video scene cuts and align generated musical chords with them. Unlike existing models, our approach retains event-based encoding, ensuring fine-grained timing control and expressive musical nuances. We also propose a mapping scheme to bridge the video emotion classifier, which produces discrete emotion categories, with the emotion-conditioned MIDI generator, which operates on continuous-valued valence-arousal inputs. Our method outperforms state-of-the-art models in objective and subjective evaluations across different video datasets, demonstrating its effectiveness in generating music aligned to video both emotionally and temporally.
1 Introduction
@@
The online distribution of user-generated multimedia content is expanding at an exponential rate thanks to affordable, high-quality recording equipment and video editing software (Falkowski-Gilski and Uhl 2020). A key challenge in content creation is providing suitable soundtracks to enhance viewer engagement (Buhler 2018). However, unauthorized use of commercially published music infringes copyright, preventing monetization for creators on platforms like YouTube. Alternatives such as purchasing music, hiring composers, or searching for royalty-free tracks are often costly, time-consuming, or fail to ensure proper synchronization with video content. Automatic video-based music generation offers a promising solution to this problem (Di et al. 2021; Zhuo et al. 2023; Kang, Poria, and Herremans 2024).
@@
Existing approaches that generate music as audio waveforms lack editability (Lam et al. 2024; Copet et al. 2024). These methods compress all stages of music production, including composition, performance, editing, mixing, and mastering, into a single process, which limits creative control for musicians and producers. By contrast, generating music in a symbolic format like MIDI (Musical Instrument Digital Interface) offers greater flexibility. MIDI functions as a digital music score, encoding instrument names, note pitches, durations, and velocities. Musicians and producers can edit compositions, synthesize audio using digital audio workstations (DAWs), or record and refine performances.
@@
This work focuses on generating music in MIDI format from arbitrary videos using deep neural networks (DNNs). We use “MIDI” and “music” interchangeably, as MIDI is our exclusive format. A major challenge in training video-to-MIDI DNNs is the absence of large-scale paired video-MIDI datasets. Existing datasets are either domain-specific, such as pianist hand videos (Koepke et al. 2020), or contain limited samples, around 1k (Di et al. 2021; Zhuo et al. 2023). To address this, we develop a model that generates music for any video type by leveraging the Lakh MIDI Dataset (Raffel 2016), the largest available collection with 176,581 samples, to ensure diverse and high-quality outputs.
@@
Since the Lakh MIDI dataset lacks corresponding videos, we adopt a two-stage approach: extracting video features relevant to music generation and using them as conditioning inputs. Drawing from musicology literature, we identify two essential aspects for effective video-based music generation: emotional alignment (Deutsch 2007) and temporal synchronization (Prendergast 1992). We name our model EMSYNC, as it aligns music with video by matching their emotions, and synchronizing their temporal boundaries. While our method is applicable to any selection of temporal boundary and emotion representation, we define musical boundaries using long-duration chords, video boundaries using scene cuts, and represent emotions with the valence-arousal model (Russell 1980). Specifically, we extract scene cut locations from the input video and guide the music generator to produce long-duration chords near those locations that are rhythmically and harmonically compatible with the rest of the generated music. We also use a pretrained video emotion classifier to estimate discrete emotion probabilities, map them to valence-arousal values, and condition our music generator accordingly.
@@
Temporal conditioning in MIDI-generating models presents unique challenges. While deep transformer models can handle time-based data, such as videos, their sequence dimension correlates linearly with time due to a fixed frame rate (Arnab et al. 2021; Liu et al. 2022). In contrast, MIDI is typically processed using an event-based representation, where sequence and time dimensions are not linearly correlated (Oore et al. 2020). In event-based MIDI encoding, two primary token types exist: “note” and “time shift”. Note tokens represent pitch, while time shift tokens advance the time axis, capturing both note durations and silences. Each time shift token specifies a time increment, forming a one-dimensional sequence where the position in the sequence does not directly correspond to the position in time. The key advantage of event-based encoding is the absence of a fixed time grid, allowing for subtle, expressive timing variations that reflect human musicianship. Unlike state-of-the-art video-based music generators that rely on fixed, coarse time grids (Di et al. 2021; Kang, Poria, and Herremans 2024), we introduce a novel temporal conditioning method that preserves event-based encoding, enabling fine-grained temporal control.
@@
The selection of temporal features for video-music synchronization is also non-trivial. Existing work matches video motion speed and motion saliency with musical note density, producing a dense representation of temporal dynamics that leads to continuous conditioning and frequent changes in note density (Di et al. 2021). Our inspection of the output samples from this approach identified two key limitations. First, the continuously varying note density can disrupt rhythmic consistency, making the underlying beat unstable and harder to follow. Second, because the mapping is dense, synchronization between music and video is difficult to perceive unless there are strong contrasts in motion speed and note density across consecutive sections. Motivated by works in musicology and soundtrack composition, we choose a temporal representation of video using sections, separated by sparse temporal boundaries of scene cuts (Cohen 1990; Prendergast 1992). To enable our music generator to anticipate future boundaries and align its output accordingly, we propose using boundary offsets. We conduct objective and subjective evaluations to measure the temporal synchronization between input video and generated music, and demonstrate that our approach outperforms dense temporal conditioning.
In addition, we tackle the challenge of using emotions as an intermediary to link music generation to video. Emotions are commonly represented with either categorical (Ekman 1971) or dimensional (continuous-valued) (Russell 1980) models. Based on prior findings from psychology and emotion research (Russell and Mehrabian 1977), we provide a mapping between categorical and dimensional emotion representations. This mapping enables the combined use of different emotion representations across data modalities, such as categorical emotions for video classification and dimensional emotions for MIDI generation, as in our case. This approach ensures flexibility and supports the use of training data labeled with either type of emotion representation.
@@
We compare our method to state-of-the-art video-based MIDI generators through objective and subjective evaluation on different datasets, where it outperforms the state-of-the-art methods. Our main contributions are as follows:
@@
  • We develop a video-based MIDI generator that surpasses state-of-the-art models across all subjective and a majority of the objective metrics for all datasets.
@@
  • We introduce boundary offsets for temporal conditioning in transformers, ensuring precise alignment between the input video and the output MIDI.
@@
  • Based on psychological research, we establish a mapping between discrete emotion categories and continuous valence-arousal values, an enabling emotional connection between the input video and the generated music.
2 Related work
@@
While our method applies to arbitrary videos, several works focus on generating symbolic music for specific types of videos, such as those featuring human movements like dancing or instrumental performances. The Foley Music model (Gan et al. 2020) generates MIDI from videos of musicians by processing body keypoint movements using a Graph Convolutional Network (Kipf and Welling 2017) and a Transformer (Vaswani et al. 2017). Similarly, Koepke et al. (2020) and Su et al. (2020) use deep neural networks with residual connections to generate symbolic music from videos of finger movements on piano keyboards. Due to their specialized nature, these approaches rely on datasets containing video-MIDI pairs, however, these datasets typically contain fewer than 1k samples (Zhao et al. 2018; Li et al. 2018; Koepke et al. 2020; Su, Liu, and Shlizerman 2020). The RhythmicNet model employs a multi-stage process to generate music from dance videos by predicting beats and style, generating a drum track, and subsequently creating multitrack music (Su, Liu, and Shlizerman 2021).
Some studies explore the more general task of generating symbolic music for arbitrary videos. The most similar to our approach, the Controllable Music Transformer (CMT), generates music based on video features such as motion speed, motion saliency, and timing (Di et al. 2021). CMT employs the Lakh Pianoroll Dataset (Dong et al. 2018), the same dataset we use, and processes music using an extended Compound Word representation, where each token encodes type, beat/bar marking, note strength, note density, instrument, pitch, and duration (Hsiao et al. 2021). While this representation reduces sequence length, it significantly increases input dimensionality. Consequently, CMT uses a low temporal resolution of four ticks per beat, translating to a time resolution of 125 milliseconds for a song at 120 beats per minute. Furthermore, note events align precisely with beat subdivisions, whereas human performance naturally deviates from beat subdivisions due to expressive timing (Fabian, Timmers, and Schubert 2014). In contrast, our model utilizes an event-based representation with independent time shift tokens at an 8 ms resolution. This resolution is sufficient to capture musical nuance while maintaining absolute time rather than rigid beat subdivisions (Oore et al. 2020). Additionally, unlike CMT, our model integrates high-level emotional conditioning alongside low-level temporal features.
@@
In a follow-up to CMT, Zhuo et al. introduced the Symbolic Music Videos dataset, containing video-MIDI pairs (2023). They sourced MIDI data from piano tutorial videos, automatically transcribing audio using the Onsets and Frames model, resulting in 1140 samples (Hawthorne et al. 2018). Video features such as color histograms, RGB frame differences for motion, and high-level CLIP features (Radford et al. 2021) were extracted to train three models sequentially: one generating chord sequences, another generating the melody, and a third generating the accompaniment.
@@
The Video2Music model shares similarities with our approach by utilizing both low-level video features and high-level emotional conditioning, but differs in application (Kang, Poria, and Herremans 2024). The authors compiled the MuVi-Sync dataset, consisting of 748 music videos labeled with musical attributes like note density, loudness, chords, and key. Their encoder-decoder transformer takes low- and high-level video features, along with user-provided primer chord and key, to generate chord sequences. These sequences are then arpeggiated using fixed patterns to create the final MIDI output. However, this method’s reliance on a fixed time grid omits the subtle expressive timing found in human performances. Additionally, its use of fixed arpeggiation patterns limits musical diversity, and requiring user-defined chord and key inputs restricts accessibility for non-musicians.
3 Methodology
@@
In this section, we present our video-based symbolic music generator. We match the output music to the input video based on emotions and temporal boundaries. Since there are no datasets where videos are paired with symbolic music, we employ a two-stage approach, i.e., use independent modules for video analysis and conditional music generation, and then combine them. Our pipeline is shown in Figure 1. We first highlight our music generator and how it is conditioned on emotions and temporal boundaries. We then describe how we incorporate temporal and emotional video features into our conditional music generator.
3.1 Conditional music generator
We extend an existing emotion-based MIDI generator trained with valence-arousal labels (Sulun, Davies, and Viana 2022) assigned to the Lakh MIDI dataset (Raffel 2016). By introducing temporal boundary conditioning and emotion mapping, we align music generation with input videos at inference time. Figure 2 illustrates the music generator, with the emotion-conditioning mechanism shown in the upper part and the boundary-conditioning mechanism in the lower part.
@@ 3.1 Conditional music generator
Figure 1: Video-based music generation pipeline. The text and the image next to the arrows demonstrate sample values.
Training dataset and preprocessing
@@
We train our music generator on the Lakh Pianoroll Dataset (LPD) (Dong et al. 2018), which contains 174,154 pianorolls derived from the Lakh MIDI Dataset (Raffel 2016). We tokenize the pianorolls using an event-based symbolic music representation (Oore et al. 2020). Specifically, an ON (note on) token marks the start of a note, and an OFF (note off) token marks its end. These tokens also encode pitch and instrument information. For example, a piano note with a MIDI pitch of 60 (C4) is denoted as PIANO_ON_60. Since a larger number of instruments increases vocabulary size, we use the Lakh Pianoroll Dataset-5 variant, where all instrument tracks are merged into five predefined categories: bass, drums, guitar, piano, and strings (Dong et al. 2018). However, our method is adaptable to datasets with different instrument groupings.
@@
TIMESHIFT tokens are used to move along the time axis, representing both note durations and the silences between them. Each token specifies a time increment in milliseconds. For example, an 800 ms shift is encoded as TIMESHIFT_800. We use a temporal resolution of 8 ms with a maximum shift of 1000 ms. Longer durations are represented using multiple consecutive TIMESHIFT tokens. We also use the START tokens to mark the beginning of the songs, the BAR tokens to indicate the musical bars, and the PAD tokens to standardize input sequence lengths in minibatches. We introduce the CHORD token to mark long-duration chords as temporal boundaries, as they often serve as anchor points in musical compositions (Bharucha and Krumhansl 1983). However, this is a design choice, and our method can accommodate any selection of temporal boundaries.
@@
To fully exploit available training data, we implement several strategies. Since we aim to generate multi-instrument compositions, we prioritize pieces with three or more instruments. However, one third of the Lakh MIDI dataset contains songs with only one or two instruments, with classical piano pieces dominating the single-instrument subset (Raffel 2016). As these samples still contain valuable information on musical structure, we avoid filtering them out. Instead, we prepend special tokens: FEWER_INSTRUMENTS for songs with two or fewer instruments and MORE_INSTRUMENTS for those with three or more. These tokens allow users to specify instrumentation preferences at inference time while leveraging the entire dataset during training. Additionally, to utilize MIDI samples without valence-arousal labels, we allow the model to accept NaN (Not A Number) values for valence or arousal, enabling both conditional and unconditional generation. If valence or arousal is unspecified, we assign it a NaN value and use a learned vector in place of its projection vector.
Figure 2: Our music generator. Numbers underneath valence and arousal are sample values.
Temporal boundary matching

### v3

@@ Abstract
Providing soundtracks for videos remains a costly and time-consuming challenge for multimedia content creators. We introduce EMSYNC, an automatic video-based symbolic music generator that creates music aligned with a video’s emotional content and temporal boundaries. It follows a two-stage framework, where a pretrained video emotion classifier extracts emotional features, and a conditional music generator produces MIDI sequences guided by both emotional and temporal cues. We introduce boundary offsets, a novel temporal conditioning mechanism that enables the model to anticipate upcoming video scene cuts and align generated musical chords with them. We also propose a mapping scheme that bridges the discrete categorical outputs of the video emotion classifier with the continuous valence-arousal inputs required by the emotion-conditioned MIDI generator, enabling seamless integration of emotion information across different representations. Our method outperforms state-of-the-art models in objective and subjective evaluations across different video datasets, demonstrating its effectiveness in generating music aligned to video both emotionally and temporally. Our demo and output samples are available at https://serkansulun.com/emsync.
I Introduction
@@
The online distribution of user-generated multimedia content is expanding at an exponential rate thanks to affordable, high-quality recording equipment and video editing software [11]. A key challenge in content creation is providing suitable soundtracks to enhance viewer engagement [4]. However, unauthorized use of commercially published music infringes copyright, preventing monetization for creators on platforms like YouTube. Alternatives such as purchasing music, hiring composers, or searching for royalty-free tracks are often costly, time-consuming, or fail to ensure proper synchronization with video content. Automatic video-based music generation offers a promising solution to this problem [56, 8, 21, 46].
@@
Existing approaches generate music as audio waveforms, which lack editability [28, 6] because they collapse composition, performance, and mastering into a single, inseparable signal. By contrast, generating music in a symbolic MIDI format offers greater flexibility by maintaining these stages as discrete, modifiable events. MIDI functions as a digital score, encoding instrument names, note pitches, durations, and velocities. Because these elements are editable, musicians can modify compositions or refine performances in digital audio workstations (DAWs).
@@
This work focuses on generating music in MIDI format from arbitrary videos using deep neural networks (DNNs). We use “MIDI” and “music” interchangeably, as MIDI is our exclusive format. A major challenge in training video-to-MIDI DNNs is the absence of large-scale paired video-MIDI datasets. Existing datasets are either domain-specific, such as pianist hand videos [26] or MIDI files with lyrics metadata [42], or contain limited samples, around 1k [8, 56]. To address this, we develop a model that generates music for any video type by leveraging the Lakh MIDI Dataset [36], the largest available collection with 176,581 samples, to ensure diverse and high-quality outputs.
@@
Since the Lakh MIDI dataset lacks corresponding videos, we adopt a two-stage approach: extracting video features relevant to music generation and using them as conditioning inputs. Drawing from the musicology literature, we identify two essential aspects for effective video-based music generation: emotional alignment [7] and temporal synchronization [34]. We name our model EMSYNC, as it aligns music with video by matching their emotions, and synchronizing their temporal boundaries. While our method is applicable to any selection of temporal boundary and emotion representation, we define musical boundaries using long-duration chords, video boundaries using scene cuts, and represent emotions with a valence-arousal model [38]. Specifically, we extract scene cut locations from the input video and guide the music generator to produce long-duration chords near locations that are rhythmically and harmonically compatible with the rest of the generated music. We also use our pretrained video emotion classifier to estimate discrete emotion probabilities [44, 43], map them to valence-arousal values, and condition our music generator accordingly.
@@
Temporal conditioning in MIDI-generating models presents unique challenges. While deep transformer models can handle time-based data, such as videos, their sequence dimension correlates linearly with time due to a fixed frame rate [1, 30]. In contrast, MIDI is processed using an event-based representation in modern applications [8, 21, 46], where sequence and time dimensions are not linearly correlated [31]. In event-based MIDI encoding, two primary token types exist: “note” and “time shift”. Note tokens represent pitch, while time shift tokens advance the time axis, capturing both note durations and silences. Each time shift token specifies a time increment, forming a one-dimensional sequence where the position in the sequence does not directly correspond to the position in time. The key advantage of event-based encoding is the absence of a fixed time grid, allowing for expressive timing variations that reflect human musicianship. Unlike state-of-the-art video-based music generators that rely on fixed, coarse time grids [56, 8, 21, 46], we introduce a novel temporal conditioning method that preserves event-based encoding, enabling fine-grained temporal control.
@@
The selection of temporal features for video-music synchronization is also non-trivial. Existing work matches video motion speed and motion saliency with musical note density, producing a dense representation of temporal dynamics that leads to continuous conditioning and frequent changes in note density [8, 46, 56, 21]. Our inspection of the output samples from this approach identified two key limitations. First, the continuously varying note density can disrupt rhythmic consistency, making the underlying beat unstable and harder to follow. Second, because the mapping is dense, synchronization between music and video is difficult to perceive unless there are strong contrasts in motion speed and note density across consecutive sections. Motivated by works in musicology and soundtrack composition, we choose a temporal representation of video using sections, separated by the sparse temporal boundaries of scene cuts [5, 34]. By focusing on these sparse boundaries rather than continuous video features, the model only needs to align with discrete temporal points instead of constantly modifying note density, therefore avoiding potential rhythmic instabilities. To achieve this, we introduce boundary offsets, which are auxiliary inputs for each token representing the time remaining until the next scene cut. This allows the autoregressive generator to anticipate upcoming boundaries at every step and align generated musical events with them, ensuring precise synchronization. We conduct objective and subjective evaluations to measure the temporal synchronization between input video and generated music, and demonstrate that our approach outperforms dense temporal conditioning.
In addition, we tackle the challenge of using emotions as an intermediary to link music generation with video. Our pretrained video emotion classifier outputs probabilities of categorical emotions [44], trained on Ekman‑6, one of the largest video-emotion datasets providing categorical labels [51]. For music, we previously labeled the Lakh MIDI dataset [36] with valence-arousal values [41], producing a training set with a sample size 90 times that of the largest emotion-labeled MIDI dataset [32] and a total duration 187 times that of the longest [19]. To bridge these different emotion representations, we introduce a mapping between categorical and dimensional emotions using results from prior user studies [37]. This enables the use of larger datasets for improved performance, such as the Ekman-6 video dataset with categorical labels [51] and our previous MIDI dataset with dimensional labels [41], while supporting training on either type of emotion annotation and maintaining flexibility.
@@
We compare our method to state-of-the-art video-based MIDI generators through objective and subjective evaluation on different datasets. Our main contributions are as follows:
@@
  • We present a new framework to train a video-based MIDI generator that surpasses state-of-the-art models across all subjective and a majority of the objective metrics for all datasets.
@@
  • We introduce the concept of boundary offsets for temporal conditioning in transformers, ensuring precise alignment between the input video and the output MIDI.
@@
  • Based on psychological research [37], we propose a mapping between discrete emotion categories and continuous valence-arousal values, enabling the integration of multimodal data labeled with different emotion representations for video-to-music generation.
II Related work
@@
While our method applies to arbitrary videos, several works focus on generating symbolic music for specific types of videos, such as those featuring human movements like dancing or instrumental performances. The Foley Music model [12] generates MIDI from videos of musicians by processing body keypoint movements using a Graph Convolutional Network [25] and a Transformer [48]. Similarly, Koepke et al. [26] and Su et al. [39] use deep neural networks with residual connections to generate symbolic music from videos of finger movements on piano keyboards. Due to their specialized nature, these approaches rely on datasets containing video-MIDI pairs, however, these datasets typically contain fewer than 1k samples [55, 29, 26, 39]. The RhythmicNet model employs a multi-stage process to generate music from dance videos by predicting beats and style, generating a drum track, and subsequently creating multitrack music [40].
Some studies explore the more general task of generating symbolic music for arbitrary videos. The Controllable Music Transformer (CMT) generates music based on video features such as motion speed, motion saliency, and timing [8]. It processes music using an extended Compound Word representation, where each token encodes type, beat/bar marking, note strength, note density, instrument, pitch, and duration [17].
@@
In a follow-up to CMT, Zhuo et al. introduced the V-MusProd model alongside the Symbolic Music Videos dataset, containing video-MIDI pairs [56]. They sourced MIDI data from piano tutorial videos, automatically transcribing audio using the Onsets and Frames model, resulting in 1140 samples [14]. Video features such as color histograms, RGB frame differences for motion, and high-level CLIP features [35] were extracted to train three modules sequentially: one generating chord sequences, another generating the melody, and a third generating the accompaniment.
@@
The Video2Music model shares similarities with our approach by utilizing both low-level video features and high-level emotional conditioning, but differs in application [21]. The authors compiled the MuVi-Sync dataset, consisting of 748 music videos labeled with musical attributes like note density, loudness, chords, and key. Their encoder-decoder transformer takes low- and high-level video features to generate chord sequences. These sequences are then arpeggiated using fixed patterns to create the final MIDI output. However, relying on fixed arpeggiation patterns limits the musical diversity of the generated pieces.
The XMusic model generates MIDI for videos by analyzing the video in terms of emotional and temporal features [46]. It uses Compound Word representation and performs dense, bar-by-bar conditioning. A “selector” module trained with human feedback chooses the best music from a batch of output samples in the end. The model is trained on a privately-collected closed-source dataset.
CMT, V-MusProd, Video2Music, and XMusic all use a quantized temporal representation (32nd notes or coarser) and rely on dense temporal conditioning, often treating each musical bar independently. In contrast, our model adopts an event-based representation with independent time-shift tokens at an 8 ms resolution. This fine-grained resolution captures subtle timing variations, reflecting how humans naturally play music, rather than imposing rigid, quantized beat subdivisions [31]. Furthermore, our use of sparse temporal conditioning in absolute time eliminates reliance on musical bars, enabling the model to handle any type of MIDI, including expressive performances that may not follow strict bar structures.
III Methodology
@@
In this section, we present our video-based symbolic music generator. Since there are no datasets where videos are paired with symbolic music, we employ a two-stage approach, i.e., using independent modules for video analysis and conditional music generation, and then combine them.
In order to make our contributions self-contained, we briefly summarize the existing components, namely, our video emotion classifier and emotion-based MIDI generator from prior work. Next, we describe how we integrate these components through a mapping scheme that bridges different emotion representation modalities, along with our novel temporal conditioning mechanism for generating MIDI that is temporally aligned with the input video.
III-A Overview of prior components
III-A1 Video emotion classification
Our video emotion classifier exploits publicly available pretrained models for a multimodal video analysis and uses the resulting pretrained features to classify the emotions of arbitrary videos [44, 43]. It employs pretrained models for automatic speech recognition (ASR), optical character recognition (OCR), facial expression classification, audio classification, and image understanding. Textual features obtained from ASR and OCR are further processed using a text sentiment classifier. Cross-attention layers [48] integrate and process these multimodal features, and a linear layer produces the final emotion probabilities. The model is trained on the Ekman-6 dataset [51], consisting of 1637 videos labeled with the six basic emotions derived from the original work of Ekman: anger, disgust, fear, joy, sadness, and surprise [10].
III-A2 Emotion-based MIDI generation
We have previously collected valence and arousal values for songs in the Lakh Pianoroll Dataset [9] and built a conditional transformer to use the resulting labeled dataset [41]. To integrate continuous-valued valence and arousal features with discrete musical note tokens, we projected valence and arousal into a vector space using separate linear layers. Musical input tokens are projected into vectors of the same dimensionality using an embedding layer. These vectors are then concatenated with the projected valence and arousal vectors along the sequence dimension and fed into the transformer body with relative global attention [18].
III-B Proposed method
We extend our previous emotion-based MIDI generator [41] by introducing temporal boundary conditioning. Figure 1 illustrates the complete music generation architecture, where the upper section depicts the emotion-conditioning mechanism and the lower section shows the boundary-conditioning mechanism. Through the proposed emotion mapping, we integrate our video emotion classifier with the MIDI generator, forming a unified video-based MIDI generation framework, as shown in Figure 2.
Figure 1: Our music generator. Numbers underneath valence and arousal are sample input values for illustration. Input tokens consist of discrete melodic events (e.g., ON, OFF, and TIMESHIFT) and special functional tokens (e.g., START, BAR, and CHORD). Boundary offsets are a sequence of scalar values, illustrated as the heights of the green bars. The plus sign indicates tensor addition. The dashed magenta boxes indicate concatenation. In sequence concatenation, the encoded valence and arousal vectors with shape (1, D) and the embedded input token sequence with shape (L, D) are concatenated along the first dimension, resulting in a tensor of shape (L+2, D). In feature concatenation, the encoded boundary offsets and positional embeddings, both with shape (L+2, D/2), are concatenated along the second dimension, producing a tensor of shape (L+2, D).
@@ III-B Proposed method
Figure 2: Video-based music generation pipeline. The text and the image next to the arrows demonstrate sample values for illustration.
III-B1 Training dataset and preprocessing
@@
We train our music generator on the Lakh Pianoroll Dataset (LPD) [9], which contains 174,154 pianorolls derived from the Lakh MIDI Dataset [36]. We tokenize the pianorolls using an event-based symbolic music representation [31]. Specifically, an ON (note on) token marks the start of a note, and an OFF (note off) token marks its end. These tokens also encode pitch and instrument information. For example, a piano note with a MIDI pitch of 60 (C4) is denoted as PIANO_ON_60. Since a larger number of instruments increases vocabulary size, we use the Lakh Pianoroll Dataset-5 variant, where all instrument tracks are merged into five predefined categories: bass, drums, guitar, piano, and strings [9]. However, our method is adaptable to datasets with different instrument groupings.
@@
TIMESHIFT tokens are used to move along the time axis, representing both note durations and the silences between them. Each token specifies a time increment in milliseconds. For example, an 800 ms shift is encoded as TIMESHIFT_800. We use a temporal resolution of 8 ms with a maximum shift of 1000 ms. Longer durations are represented using multiple consecutive TIMESHIFT tokens. We also use the START tokens to mark the beginning of the songs, the BAR tokens to indicate the musical bars, and the PAD tokens to standardize input sequence lengths in minibatches. We introduce the CHORD token to mark long-duration chords as temporal boundaries, as they often serve as anchor points in musical compositions [2]. However, this is a design choice, and our method can accommodate any selection of temporal boundaries such as those separating melodic phrases or song sections.
@@
To fully exploit the available training data, we implement several strategies. Since we aim to generate multi-instrument compositions, we prioritize pieces with three or more instruments. However, one third of the Lakh MIDI dataset contains songs with only one or two instruments, with classical piano pieces dominating the single-instrument subset [36]. As these samples still contain valuable information on musical structure, we avoid filtering them out. Instead, we prepend special tokens: FEWER_INSTRUMENTS for songs with two or fewer instruments and MORE_INSTRUMENTS for those with three or more. These tokens allow users to specify instrumentation preferences at inference time while leveraging the entire dataset during training. Additionally, to utilize MIDI samples without valence-arousal labels, we allow the model to accept NaN (Not A Number) values for valence or arousal, enabling both conditional and unconditional generation. If valence or arousal is unspecified, we assign it a NaN value and use a learned vector in place of its projection vector.
III-B2 Temporal boundary matching